# Q-Insight: Understanding Image Quality via Visual Reinforcement Learning

### v1

@@ 4.1 Experimental Setup
Datasets and Metrics. For the score regression task, we use diverse IQA datasets across four categories: (a) in-the-wild datasets, including KonIQ [15], SPAQ [10], and LIVE-Wild [11]; (b) synthetic distortion datasets, including KADID [21] and CSIQ [19]; (c) model-processed distortions, including PIPAL [13]; and (d) AI-generated images from AGIQA [20]. Following [56], we split KonIQ into training and test sets, with approximately $7000$ training images. Mean Opinion Scores (MOS) across these datasets are normalized into the range $[1,5]$. The remaining datasets are exclusively used to evaluate the model’s out-of-distribution (OOD) generalization capability. For the degradation perception task, we randomly select $7000$ images from DQ-495K [57] that contain a single distortion for training, with an additional $1000$ images reserved for testing. We adopt the pearson linear correlation coefficient (PLCC) and spearman rank-order correlation coefficient (SRCC) as metrics to evaluate performance on the score regression task, following [18, 51, 56]. For the degradation perception task, we employ distortion class accuracy and degradation level accuracy as evaluation metrics.
Implementation Details. We adopt Qwen-2.5-VL-7B [1] as our base model. In the GRPO algorithm, the generation number $N$ is set to 8, while the weights $\alpha_{1}$ and $\alpha_{2}$ are set to $0.25$ and $0.75$, respectively. The threshold $\epsilon$ is set to $0.35$. We employ AdamW [25] as the optimizer, using an initial learning rate of $1\times 10^{-6}$ that linearly decays during training. The model is trained for 10 epochs with a total batch size of 128. Training is completed in approximately one day using 16 NVIDIA A100 GPUs.
@@ 4.2 Score Regression
We first evaluate our Q-Insight on the score regression task. We compare Q-Insight with handcrafted methods NIQE [28] and BRISQUE [27]; non-MLLM deep-learning methods including NIMA [44], HyperIQA [41], DBCNN [65], MUSIQ [18], CLIP-IQA+ [46], and ManIQA [54]; and recent MLLM-based methods such as C2Score [73], Q-Align [51], DeQA-Score [56], and a supervised fine-tuned Qwen [1]. For a fair comparison, all methods (except handcrafted ones) are trained on the KonIQ dataset, and all MLLM-based methods utilize approximately 7B parameters. The comparison results in terms of PLCC and SRCC between Q-Insight and other IQA methods are presented in Tab. 3.4. Compared with the state-of-the-art method DeQA-Score, our Q-Insight performs slightly worse on the in-domain KonIQ dataset. However, on out-of-distribution (OOD) datasets, Q-Insight consistently outperforms all baseline methods across nearly all benchmarks, achieving approximately $0.2$ improvements in both PLCC and SRCC. This demonstrates the effectiveness and strong generalization capability of our approach. Fig. 3 shows two illustrative cases of the reasoning process in the score regression task. Notably, our Q-Insight does not merely output a numerical score; it also provides a comprehensive and detailed reasoning process. In the first example (top of Fig. 3), Q-Insight successfully recognizes the text on the neon sign and carefully analyzes details such as lighting conditions. Similarly, the second example (bottom of Fig. 3) highlights Q-Insight’s capability in understanding image composition. This further demonstrates that Q-Insight moves beyond merely regressing scores, critically analyzing image quality issues from multiple perspectives, and advancing toward a more comprehensive understanding of image quality.
Table 4: Ablation study on the score regression task between multi-task and single-task training. Ours with joint-training significantly outperforms score-only training on PLCC / SRCC metrics.
@@
Ours 0.918 0.903 0.702 0.458 0.870 0.816 0.685 0.765
(Score-Only) / 0.895 / 0.899 / 0.702 / 0.435 / 0.839 / 0.766 / 0.640 / 0.739
Ours 0.933 0.907 0.742 0.486 0.893 0.811 0.870 0.806
(Joint-Training) / 0.916 / 0.905 / 0.736 / 0.474 / 0.865 / 0.764 / 0.824 / 0.783
@@ 4.3 Distortion Perception
Figure 4: Subjective ablation comparison between joint multi-task training and w/o joint training on the explanation of image scoring. With joint training, our method can better perceive degradation cues in images (such as pixelated appearance), thereby improving the accuracy of quality assessment.
Figure 5: Image comparison reasoning results of our Q-Insight and DepictQA [58, 57]. Q-Insight outperforms DepictQA in comprehensive content understanding and accurate degradation perception.
@@ 4.4 Ablation Studies
Effect of multi-task training. To demonstrate the effectiveness of multi-task training, we compare our jointly trained Q-Insight with two single-task variants, trained separately on each task. The comparison results are presented in Tabs. 4.2 and 4.2. As shown in Tab. 4.2, the jointly trained Q-Insight significantly outperforms the score-only variant on nearly all datasets, particularly those with synthetic distortions (KADID, CSIQ) and model-processed distortions (PIPAL). This clearly indicates that incorporating the degradation perception task benefits the score regression task. Fig. 4 further illustrates the importance of multi-task training: Q-Insight can better perceive detailed degradations such as pixel-level artifacts, thus enhancing overall accuracy in quality assessment. Similarly, Tab. 4.2 demonstrates that in the degradation perception task, the jointly trained model outperforms the degradation-only variant across all distortion categories. This suggests that the score regression task also positively contributes to the degradation perception capability. These results confirm the mutual benefit and effectiveness of our multi-task training approach. Furthermore, this clearly demonstrates that the visual quality understanding potential of MLLMs can be significantly enhanced through carefully designed training tasks and objectives.
4.5 Zero-Shot Image Comparison Reasoning
@@
Our Q-Insight effectively generalizes to zero-shot image comparison reasoning tasks, covering both reference-based and non-reference-based scenarios, as illustrated in Figs. 1 and 5. In Fig. 1, Q-Insight conducts a reference-based comparison, where the reference image is of lower quality, and Images A and B are outputs from different super-resolution methods. Q-Insight compares these two images from both fidelity and overall quality perspectives, demonstrating its potential utility in image enhancement tasks. In Fig. 5, Q-Insight surpasses DepictQA [58, 57] in terms of comprehensive content understanding and accurate degradation perception. Importantly, Q-Insight achieves this without relying on specialized training on large-scale textual datasets, highlighting its strong generalization ability enabled by the GRPO-based framework and multi-task training strategy.
@@ 5 Conclusion
In this paper, we introduce Q-Insight, a novel GRPO-based model for comprehensive image quality understanding. It jointly optimizes score regression and degradation perception tasks using only a limited amount of labeled data. Unlike traditional methods that rely on extensive textual annotations or purely numerical scoring, our framework combines numerical accuracy with interpretative reasoning, significantly improving the perceptual analysis capabilities of image quality models. Extensive experiments show that Q-Insight consistently outperforms existing state-of-the-art methods across various datasets and tasks, demonstrating impressive zero-shot generalization to image comparison reasoning. Looking ahead, Q-Insight can extend its capabilities to a wide range of tasks, such as image aesthetic assessments, and serve as a powerful discriminative signal to improve image enhancement models. As a unified model for scoring, perception, comparison, and reasoning, Q-Insight can act as a central hub, coordinating image reconstruction tools and providing valuable insights into the enhancement process. This integrated and automated system has the potential to revolutionize image quality understanding and enhancement, providing a unified solution that can transform how image quality is evaluated, improved, and applied across various fields.

### v2

@@ 4.1 Experimental Setup
Datasets and Metrics. For the score regression task, we use diverse IQA datasets across four categories: (a) in-the-wild datasets, including KonIQ [16], SPAQ [11], and LIVE-Wild [12]; (b) synthetic distortion datasets, including KADID [22] and CSIQ [20]; (c) model-processed distortions, including PIPAL [14]; and (d) AI-generated images from AGIQA [21]. Following [58], we split KonIQ into training and test sets, with approximately $7000$ training images. Mean Opinion Scores (MOS) across these datasets are normalized into the range $[1,5]$. The remaining datasets are exclusively used to evaluate the model’s out-of-distribution (OOD) generalization capability. For degradation perception task, we randomly select $7000$ images from DQ-495K [59] that contain a single distortion for training, with an additional $1000$ images reserved for testing. We adopt the pearson linear correlation coefficient (PLCC) and spearman rank-order correlation coefficient (SRCC) as metrics to evaluate performance on score regression task, following [19, 53, 58]. For degradation perception, we use the accuracy of distortion class and degradation level as evaluation metrics.
Implementation Details. We adopt Qwen-2.5-VL-7B-Instruct [1] as our base model. In the GRPO algorithm, the generation number $N$ is set to $8$, the weight of KL divergence penalty $\beta$ is set to $1\times 10^{-3}$, while the weights $\alpha_{1}$ and $\alpha_{2}$ are set to $0.25$ and $0.75$, respectively. The threshold $\epsilon$ is set to $0.35$. We employ AdamW [26] as the optimizer, using an initial learning rate of $1\times 10^{-6}$ that linearly decays to $1\times 10^{-9}$ during training. The model is trained for $10$ epochs with a total batch size of $128$. Training is completed in approximately one day using 16 NVIDIA A100 GPUs.
@@ 4.2 Score Regression
We first evaluate our Q-Insight on the score regression task. We compare Q-Insight with handcrafted methods NIQE [29] and BRISQUE [28]; non-MLLM deep-learning methods including NIMA [47], MUSIQ [19], CLIP-IQA+ [49], and ManIQA [56]; and recent MLLM-based methods such as C2Score [74], Q-Align [53], DeQA-Score [58], and a supervised fine-tuned Qwen [1]. For a fair comparison, all methods (except handcrafted ones) are trained on the KonIQ dataset, and all MLLM-based methods utilize approximately 7B parameters. The comparison results in terms of PLCC and SRCC between Q-Insight and other IQA methods are presented in Tab. 3.4. Compared with the state-of-the-art method DeQA-Score, our Q-Insight performs slightly worse on the in-domain KonIQ dataset. However, on out-of-distribution (OOD) datasets, Q-Insight consistently outperforms all baseline methods across nearly all benchmarks, achieving approximately $0.02$ improvements in both PLCC and SRCC. This demonstrates the effectiveness and strong generalization capability of our approach. Fig. 3 illustrates two cases showing the reasoning capability in the score regression task. Specifically, our method goes beyond merely outputting numerical scores and provides detailed, structured reasoning. In the first case (top of Fig. 3), Q-Insight correctly identifies and analyzes textual information displayed on a neon sign, thoroughly examining details such as lighting conditions. In the second case (bottom of Fig. 3), Q-Insight demonstrates its strength in interpreting image composition aspects, such as the arrangement of visual elements and the primary focal point of the image. These examples further illustrate how Q-Insight advances beyond score regression task, offering valuable insights into image quality by examining various perceptual factors from multiple perspectives, ultimately contributing to a comprehensive understanding of image quality.
Table 3: Ablation study on the score regression task between multi-task and single-task training. Q-Insight with joint-training significantly outperforms score-only training on PLCC / SRCC metrics.
@@
Ours (Score-Only) 0.918 0.903 0.702 0.458 0.870 0.816 0.685 0.765
/ 0.895 / 0.899 / 0.702 / 0.435 / 0.839 / 0.766 / 0.640 / 0.739
Ours (Joint-Training) 0.933 0.907 0.742 0.486 0.893 0.811 0.870 0.806
/ 0.916 / 0.905 / 0.736 / 0.474 / 0.865 / 0.764 / 0.824 / 0.783
@@ 4.3 Distortion Perception
Figure 4: Image comparison reasoning results of our Q-Insight and DepictQA [60]. Q-Insight outperforms DepictQA in comprehensive content understanding and accurate degradation perception.
@@ 4.4 Ablation Studies
Effect of multi-task training. To validate the effectiveness of multi-task training, we compare our jointly trained Q-Insight model with two single-task variants, each trained independently on a single task. The comparison results are presented in Tabs. 4.2 and 4.2. As shown in Tab. 4.2, the jointly trained Q-Insight significantly outperforms the score-only variant on nearly all datasets, especially on datasets involving synthetic distortions (KADID [22], CSIQ [20]) and those containing model-generated distortions (PIPAL [14]). This demonstrates that incorporating the degradation perception task can effectively enhance performance in the score regression task. Fig. A in the appendix further presents the benefits of multi-task training, showing that Q-Insight can precisely identify detailed degradations such as pixel-level artifacts, thereby improving overall accuracy in quality assessment. Similarly, Tab. 4.2 indicates that in the degradation perception task, our jointly trained model consistently surpasses the degradation-only variant across all distortion types. This suggests that the score regression task also positively contributes to degradation perception capabilities. These experimental results verify the mutual benefit and effectiveness of the proposed multi-task training strategy. Moreover, our findings clearly show that the visual quality understanding potential of MLLMs can be significantly improved through carefully designed training tasks and learning objectives.
Ablation on the score threshold $\epsilon$. Introducing the threshold allows the model’s predictions to vary within an acceptable margin. Tab. B.1 in the appendix reports the ablation results for different choices of $\epsilon$. Q-Insight consistently achieves robust and stable performance across various threshold values, demonstrating that its effectiveness does not depend on careful tuning of $\epsilon$.
Table 5: Accuracy and PLCC / SRCC results of the reference-based comparison task on the SRbench [6]. Reg-Acc and Gen-Acc represent the accuracy between regression-based and generation-based restoration methods, respectively. Q-Insight outperforms score- and description-based methods.
{NiceTabular}
[…]
Score-Based PSNR 80.07% 41.70% 34.70% – –
SSIM [50] (TIP 04) 83.00% 45.30% 37.40% – –
LPIPS [66] (CVPR 18) 82.00% 63.90% 65.80% – –
A-FINE [6] (CVPR 25) 83.30% 78.90% 82.40% – –
Description-Based DepictQA [60] (ECCV 24) 73.00% 61.64% 62.96% 0.3457 0.3412
Q-Insight (Zero-Shot) 78.67% 68.64% 75.51% 0.6385 0.6297
Q-Insight (Trained) 85.67% 77.78% 82.80% 0.7627 0.7614
4.5 Image Comparison Reasoning
@@
Our Q-Insight effectively generalizes to zero-shot image comparison reasoning tasks in both reference-based and non-reference-based scenarios, as illustrated in Figs. 1 and 4. Specifically, Fig. 1 shows a reference-based comparison scenario, where the reference image is of lower quality, and Images A and B are outputs generated by two different super-resolution methods. Fig. 4 demonstrates Q-Insight’s superiority over DepictQA [60], highlighting its enhanced content understanding and precise perception of degradations. These examples illustrate Q-Insight’s robust generalization ability, largely enabled by its RL-based framework and multi-task training strategy.
Furthermore, the comparison reasoning performance can be further boosted by training on a small number of labeled comparison pairs using reinforcement learning. Specifically, we randomly sample 5k data pairs from the DiffIQA [6] dataset, where each pair is labeled only with comparison results, without any textual descriptions. Results are shown in Tab. 4.4. Our Q-Insight consistently surpasses all score-based and description-based methods in terms of overall accuracy and PLCC/SRCC metrics, demonstrating its promising applicability to various image enhancement tasks. Notably, even the zero-shot version of Q-Insight substantially outperforms DepictQA [60], despite the latter relying on large-scale textual datasets. Additionally, A-Fine [6] utilizes more than 200k data pairs collected from four different datasets, combined with a complex three-stage training pipeline, thus requiring over 40 times more data and considerable effort to develop optimal training strategies. In contrast, Q-Insight achieves superior performance through relatively straightforward yet highly effective visual reinforcement learning. Further details and additional results are provided in Sec. C of the Appendix.
@@ 5 Conclusion
In this paper, we introduce Q-Insight, a novel GRPO-based model for comprehensive image quality understanding. It jointly optimizes score regression and degradation perception tasks using only a limited amount of labeled data. Unlike traditional methods that rely on extensive textual annotations or purely numerical scoring, our framework combines numerical accuracy with interpretative reasoning, significantly improving the perceptual analysis capabilities of image quality models. Extensive experiments show that Q-Insight consistently outperforms existing state-of-the-art methods across various datasets and tasks, demonstrating impressive zero-shot generalization and superior comparison reasoning capability. Looking ahead, Q-Insight can extend its capabilities to a wide range of tasks, such as image aesthetic assessments, and serve as a powerful discriminative signal to improve image enhancement models. As a unified model for scoring, perception, comparison, and reasoning, Q-Insight can act as a central hub, coordinating image reconstruction tools and providing valuable insights into the enhancement process. This integrated and automated system has the potential to revolutionize image quality understanding and enhancement, providing a unified solution that can transform how image quality is evaluated, improved, and applied across various fields.
Limitations. While achieving promising performance, Q-Insight focuses primarily on natural images. Extending to AI-generated images and videos remains essential and is reserved for future exploration.
Appendix A Designed Prompts
The prompts designed for each task are detailed in Tab. A. Specifically, a general system prompt is shared across tasks, which encourages the model to explicitly output its reasoning process and provide structured responses. This general prompt is supplemented by task-specific prompts tailored for score regression and degradation perception, respectively. For the score regression task, the input includes a task-specific prompt and the image to be rated, with the Mean Opinion Score (MOS) serving as the ground-truth. In the degradation perception task, the input consists of a prompt and an image characterized by a specific distortion class and severity level. We define five distortion categories: “noise”, “blur”, “JPEG”, “darken”, and “null”, where “null” indicates no distortion. Each distortion type has five severity levels: “slight”, “moderate”, “obvious”, “serious”, and “catastrophic”. The distortion class and corresponding severity level constitute the ground-truth labels. Additionally, for the comparative reasoning scenario, the inputs include a prompt, two images to be compared, and an optional reference image.
Table A: Prompts for Different Tasks. The system prompt is shared across all tasks, while task-specific prompts are additionally designed for each individual task.
Prompt for Non-Reference-Based Image Comparison: Given Image A: <image_A> and <Image_B>, assess the visual quality of both the Image A and the Image B, explain and justify which one is better considering composition and degradation. Your answer should be "Image A" or "Image B".
Prompt for Reference-Based Comparison Reasoning: Given a low-quality reference image and two enhanced outputs. Reference Image: <ref_image>, Image A: <image_A> and Image B: <image_B>. Decide which enhanced image is superior or if they are comparable. Evaluate based on: 1) fidelity and consistency with the reference image; 2) overall perceptual quality. Return exactly one of: "Image A", "Image B", or "Similar".
Appendix B More Results of Ablation Studies
B.1 Ablation study of the threshold $\epsilon$
When the difference between the model’s predicted score and the ground-truth score is less than $\epsilon$, the model receives the score reward, enabling predictions to fluctuate within an acceptable range. The ablation study on the threshold $\epsilon$ is presented in Tab. B.1 in the following. Q-Insight consistently achieves robust and stable performance across various choices of $\epsilon$, demonstrating that its effectiveness does not rely on meticulous tuning of the threshold.
Table B: Ablation study on the threshold $\epsilon$ for the score regression task. Q-Insight demonstrates robust and stable performance, indicating it does not require careful tuning of the threshold $\epsilon$.
{NiceTabular}
[…]
$\epsilon=0.15$ 0.918/0.900 0.903/0.903 0.749/0.744 0.443/0.441 0.876/0.847 0.835/0.764 0.857/0.793
$\epsilon=0.25$ 0.929/0.911 0.902/0.902 0.750/0.752 0.472/0.471 0.884/0.855 0.823/0.765 0.889/0.844
$\epsilon=0.30$ 0.932/0.912 0.906/0.903 0.742/0.738 0.471/0.460 0.887/0.856 0.828/0.773 0.868/0.822
$\epsilon=0.40$ 0.930/0.912 0.904/0.901 0.721/0.723 0.446/0.428 0.880/0.853 0.819/0.769 0.873/0.831
$\epsilon=0.50$ 0.928/0.906 0.906/0.902 0.750/0.746 0.475/0.472 0.885/0.854 0.821/0.768 0.894/0.850
$\epsilon=0.35$ (Ours) 0.933/0.916 0.907/0.905 0.742/0.736 0.486/0.474 0.893/0.865 0.811/0.764 0.870/0.824
B.2 Effect of multi-task training
In the main paper, we have quantitatively demonstrated the mutual benefits derived from multi-task training. Here, we further illustrate the effectiveness of our multi-task approach qualitatively in Fig. A. Specifically, multi-task training enables Q-Insight to better perceive detailed degradations such as pixel-level artifacts, thereby improving its overall accuracy in image quality assessment.
Figure A: Subjective ablation comparison between joint multi-task training and w/o joint training on the explanation of image scoring. With joint training, our method can better perceive degradation cues in images (such as pixelated appearance), thereby improving the accuracy of quality assessment.
Appendix C Experimental Setups and More Results of Comparison Reasoning
C.1 Datasets
For training, we use the DiffIQA [6] dataset, which contains approximately 180k reference–test image pairs generated by applying diffusion-based enhancement methods to reference images of varying quality. Each reference image is paired with multiple test images, and human annotators provide preference labels through triplet-based comparisons, resulting in roughly 180k comparison pairs. Notably, A-Fine [6] aggregates datasets from DiffIQA, TID2013 [35], KADID [22], and PIPAL [14], totaling over 200k comparison pairs for training. In contrast, we randomly sample only 5k pairs from DiffIQA for our training.
For evaluation, we adopt SRIQA-Bench [6], a benchmark specifically designed to evaluate the generalization ability of IQA models in real-world super-resolution (SR) scenarios. This dataset includes 100 low-resolution (LR) reference images, each enhanced by 10 distinct SR methods encompassing both regression-based and generative models. Human raters perform exhaustive pairwise comparisons within each image group, with each comparison annotated by at least 10 annotators to ensure reliability. Since no ground-truth high-resolution reference images are provided, models must evaluate perceptual quality based solely on the degraded LR images. Thus, SRIQA-Bench poses a challenging scenario to rigorously assess the robustness of full-reference IQA models under imperfect reference conditions. we report Reg-Acc, Gen acc, and overall ACC, which denote pairwise ranking accuracy on regression-based SR methods, generation-based SR methods, and all SR outputs, respectively. These metrics measure alignment between model predictions and human judgments under varying super-resolution styles and distortion characteristics. We also report the PLCC and SRCC of Description-based methods.
C.2 Reward Design and Training Details
In the image comparison task, the model is expected to determine which image is superior or if they are comparable. Denote the predicted score of the $i$-th response as $\text{res}^{(i)}_{\text{pred}}$, we design a comparison reward function $R_{\text{comp}}$ as follows. The reward value $r^{(i)}_{\text{comp}}$ for the $i$-th response is determined by :
$r^{(i)}_{\text{comp}}=1\quad\text{if}\ \text{res}^{(i)}_{\text{pred}}=\text{res}_{\text{comp}},~{}\text{otherwise }0.$ (7)
Finally, the overall reward of $i$-th response is calculated as:
$r^{(i)}=r^{(i)}_{\text{fmt}}+r^{(i)}_{\text{comp}},$ (8)
where $r^{(i)}_{\text{fmt}}$ is the format reward. For training, we employ AdamW [26] as the optimizer, using an initial learning rate of $1\times 10^{-6}$ that linearly decays to $1\times 10^{-9}$ during training. The model is trained for $8$ epochs with a total batch size of $32$. Training is completed in approximately $20$ hours using 16 NVIDIA A100 GPUs.
C.3 More Qualitative Results
We provide quantitative results in Tab. 4.4 in the main paper. Here we present additional qualitative results of reference-based image comparison reasoning in Figs. BCD, and E. It can be observed that, despite training on only 5k comparison pairs, Q-Insight surpasses DepictQA [60] by more effectively recognizing subtle image details and distortions, thereby producing more accurate outcomes.
Figure B: Reference-based image comparison reasoning results of our Q-Insight and DepictQA [60] on image “online4” from SRBench [6].
Figure C: Reference-based image comparison reasoning results of our Q-Insight and DepictQA [60] on image “online23” from SRBench [6].
Figure D: Reference-based image comparison reasoning results of our Q-Insight and DepictQA [60] on image “realworld33” from SRBench [6].
Figure E: Reference-based image comparison reasoning results of our Q-Insight and DepictQA [60] on image “srtest55” from SRBench [6].